HETAV PATEL

Grade 11,

differentiate the following function with respect to x (tan x)^1/x

differentiate the following function with respect to x
(tan x)^1/x

Grade:12

1 Answers

Rinkoo Gupta
9 years ago

y=tanx1/x

taking log on both sides

logy=1/x.logtanx

Now diff. w. r. to x , we get

(1/y) dy/dx=(1/x).(1/tanx).(sec2x)+logtanx(-1/x2)

dy/dx=tanx1/x[1/(xsinxcosx ) – ( logtanx)/x2
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty